# Quantum metrology through spectral measurements in quantum optics

### v1

@@ IV Harnessing photon-photon correlations in quantum parameter estimation
Figure 5: Frequency-resolved Fisher information for two sensor to estimate $\theta=\Delta$. Panels (a-c) correspond to different sensor linewidths $\Gamma=\{10^{-1},1,10\}\gamma$. Each column shares the same structure. Top row: $F_{\Delta}$ (blue solid curve), $F_{\Delta}^{I}$ (blue dashed curve), $I_{\Delta}$ (black solid curve) and $g^{(2)}_{\Gamma}$ (red solid curve) in terms of one sensor-laser detuning $\Delta_{1}$, with the second fixed at $\Delta_{2}=2\Omega$. Middle row: $g^{(2)}_{\Gamma}$ in the frequency domain ($\Delta_{1},\Delta_{2}$). Bottom row: $F_{\Delta}$ (top half) and $F_{\Delta}/F_{\Delta}^{I}$ (bottom half) in the frequency domain ($\Delta_{1},\Delta_{2}$). Parameters: (a-c) $\Omega=10\gamma,\ \Delta=0,\ \varepsilon=0.5.$
The emission of a coherently driven two-level system, particularly in the Mollow regime, reveals genuine quantum correlations—evidence of non-classical light—that remain hidden in color-blind measurements but emerge through frequency-resolved detection [42, 47, 50, 45, 51, 52, 53, 40]. A central quantity that reflects such behavior is the photon-photon correlations of the filtered emission by the zero-delay two-photon cross correlation function, which can be easily computed by means of the sensor method [42, 43, 44],
@@
This quantity provides evidences of multi-photon processes, such as the so-called leapfrog processes in the Mollow ladder [49], illustrated in the inset of the middle panel in Fig. 5(a). In the case of a two-photon leapfrog process, energy conservation sets the frequency of the two emitted photons, $\Delta_{1}$ and $\Delta_{2}$ [49, 51]. There are given by the conditions: (i) $\Delta_{1}+\Delta_{2}=-\omega_{S}$ for the transition $|-\rangle\rightrightarrows|+\rangle$, (ii) $\Delta_{1}+\Delta_{2}=0$ for the transition $|\pm\rangle\rightrightarrows|\pm\rangle$, and (iii) $\Delta_{1}+\Delta_{2}=\omega_{S}$ for the transition $|+\rangle\rightrightarrows|$-$\rangle$. These two-photon transitions appear as antidiagonal lines in the insets of the second row in Fig. 5, corresponding to superbunched emission (red regions). The second row of Fig. 5 shows $g_{\Gamma}^{(2)}$ in frequency space $(\Delta_{1},\Delta_{2})$ for different sensor linewidths, $\Gamma=\{10^{-1},1,10\}\gamma$. As previously reported in Ref. [49]: (i) narrow-band filters ($\Gamma\to 0$) tend to suppress correlations, yielding mostly uncorrelated emission (white); (ii) broadband filters ($\Gamma\to\infty$) blur spectral information—color-blind measurements—, recovering standard antibunching from a two-level system; (iii) intermediate linewidths ($\Gamma\sim\gamma$) reveal multi-photon structures, capturing rich quantum correlations. These features are more cleary observed in Fig. 6(a,b).
Similarly to Sec. III [c.f. Fig. 2(c)], the two-sensor CFI, $F_{\Delta}$, inherits a non-trivial spectral structure, shown in the upper-half of the maps shown in the third row in Fig. 5. However, as was the case with the single-sensor CFI and the fluorescence spectrum, it does not directly mirror $g^{(2)}_{\Gamma}$, due to the inherently nonlinear dependence of $F_{\theta}$ on higher-order field moments. The joint probability distribution $p(n_{1},n_{2}|\theta)$ is given by a sum of nonlinear functions of the normally ordered moments of both sensors (see Appendix H),
@@
where we have defined $\hat{D}(\alpha_{1},\alpha_{2})\equiv\hat{D}(\alpha_{1})\hat{D}(\alpha_{2})$, with $\hat{D}(\alpha_{i})=\exp[(\alpha_{i}\hat{\xi}^{\dagger}_{i}-\alpha_{i}^{*}\hat{\xi}_{i})]$, since now each sensor can be independently displaced. While a direct mapping between $g^{(2)}\Gamma$ and $F_{\theta}$ is not possible, the two-photon correlation function still offers physical insight into the structure of quantum correlations relevant for metrology.
@@
Physically, this situation may occur when, for example, one arm of the HBT setup [see Fig. 1(a)] is sufficiently long that time delays between detections erase any information about photon-photon correlations.
In Fig. 5, the top panels show $F_{\Delta}^{I}$ (blue dashed) as a function of $\Delta_{1}$ for a fixed $\Delta_{2}=2\Omega$, while the lower-half panels in the third row display the ratio $F_{\theta}/F_{\theta}^{I}$, which quantifies the metrological gain from retaining cross-correlations compared to independent measurements. When $F_{\theta}/F_{\theta}^{I}>1$ (red regions), we have a metrological advantage, reaching up to five orders of magnitude near leapfrog resonances, as it is shown in the top panels via the blue dashed curves ($F_{\Delta}^{I}$). For instance, in the aforementioned region, $(\Delta_{1},\Delta_{2})\approx(-2\Omega+\delta\omega,+2\Omega)$, we find $F_{\theta}\approx 10^{5}F_{\theta}^{I}$. Generally, retaining cross-correlations ($F_{\Delta}$) outperforms the independent measurements ($F_{\Delta}^{I}$), as long as spectral resolution is guaranteed. In the case of broadband filtering, the spectral information is washed out, resulting in $F_{\Delta}\approx F_{\Delta}^{I}$ (white regions).
These features are further illustrated in Fig. 6. In panel (a), both $F_{\Delta}$ (blue solid) and $F_{\Delta}^{I}$ (blue dashed) are plotted as a function of $\Gamma$ at a representative leapfrog point $(\Delta_{1},\Delta_{2})=(-2\Omega+\delta\omega,2\Omega)$, while panel (c) displays the ratio $F_{\Delta}/F_{\Delta}^{I}$ across $\Delta_{1}$ and $\Gamma$, clearly highlighting the spectral window where quantum correlations yield maximal metrological benefit.
Figure 6: Impact of sensor linewidth $\Gamma$ on the performance of the estimation protocol for two sensors. (a) $I_{\Delta}$ (black solid), $F_{\Delta}$ (blue solid), $F_{\Delta}^{I}$ (blue dashed), and $g^{(2)}_{\Gamma}$ in terms of $\Gamma$, for a fixed value of the sensors $(\Delta_{1},\Delta_{2})=(-2\Omega+\delta\omega,2\Omega)$. The blue shaded region represents the optimal (b) $g^{(2)}_{\Gamma}$ and (c) $F_{\Delta}/F_{\Delta}^{I}$ in terms of the frequency of one sensor $\Delta_{1}$ and the sensor linewidth $\Gamma$, while having fixed $\Delta_{2}=2\Omega$. Parameters: $\Delta=0,\ \Omega=10\gamma,\ \Delta_{2}=2\Omega,\ \delta\omega=0.5\Omega,\ \varepsilon=0.5$.
@@
V Conclusions
We have presented a framework of quantum metrology based on frequency-resolved detection, enabling the exploitation of the spectral structure of radiation emitted by open quantum systems for parameter estimation. By analyzing the classical Fisher information (CFI) across sensor frequencies and linewidths, we revealed the fundamental role of spectral resolution in determining metrological performance. In particular, we identified an optimal filtering regime—where neither spectral nor temporal information is lost—that maximizes sensitivity.

### v2

@@ IV Harnessing photon-photon correlations in quantum parameter estimation
Figure 5: Frequency-resolved Fisher information for two sensors to estimate $\theta=\Delta$. Panels (a-c) correspond to different sensor linewidths $\Gamma=\{10^{-1},1,10\}\gamma$. Each column shares the same structure. Top row: $I_{\Delta}$ (grey solid), $F_{\Delta}$ (blue solid), $F_{\Delta}^{I}$ (blue dot-dashed), $F_{\Delta}^{\vec{\alpha}_{\text{opt}}}$ (blue dashed), and $g^{(2)}_{\Gamma}$ (red solid) in terms of one sensor-laser detuning $\Delta_{1}$, with the second fixed at $\Delta_{2}=2\Omega$. Middle row: $g^{(2)}_{\Gamma}$ in the frequency domain ($\Delta_{1},\Delta_{2}$). Bottom row: $F_{\Delta}$ (upper-half) and $F_{\Delta}/F_{\Delta}^{I}$ (lower-half) in the frequency domain ($\Delta_{1},\Delta_{2}$). Parameters: (a-c) $\Omega=10\gamma,\ \Delta=0,\ \varepsilon=0.5.$
The emission of a coherently driven two-level system, particularly in the Mollow regime, reveals genuine quantum correlations—evidence of non-classical light—that remain hidden in color-blind measurements but emerge through frequency-resolved detection [44, 50, 55, 51, 56, 57, 58, 41]. A central quantity that reflects such behaviour is the photon-photon correlations of the filtered emission by the zero-delay two-photon cross correlation function, which can be easily computed by means of the sensor method [44, 45, 46],
@@
This quantity provides evidences of multi-photon processes, such as the so-called leapfrog processes in the Mollow ladder [54], illustrated in the inset of the middle panel in Fig. 5(a). In the case of a two-photon leapfrog process, energy conservation sets the frequency of the two emitted photons, $\Delta_{1}$ and $\Delta_{2}$ [54, 56]. They are determined by the conditions: (i) $\Delta_{1}+\Delta_{2}=-\omega_{S}$ for the transition $|-\rangle\rightrightarrows|+\rangle$, (ii) $\Delta_{1}+\Delta_{2}=0$ for the transition $|\pm\rangle\rightrightarrows|\pm\rangle$, and (iii) $\Delta_{1}+\Delta_{2}=\omega_{S}$ for the transition $|+\rangle\rightrightarrows|$-$\rangle$. These two-photon transitions appear as antidiagonal lines in the insets of the second row in Fig. 5, corresponding to superbunched emission (red regions). The second row of Fig. 5 shows $g_{\Gamma}^{(2)}$ in frequency space $(\Delta_{1},\Delta_{2})$ for different sensor linewidths, $\Gamma=\{10^{-1},1,10\}\gamma$. As previously reported in Ref. [54]: (i) narrow-band filters ($\Gamma\to 0$) tend to suppress correlations, yielding mostly uncorrelated emission (white); (ii) intermediate linewidths ($\Gamma\sim\gamma$) reveal multi-photon structures, capturing rich quantum correlations; (iii) broadband filters ($\Gamma\to\infty$) blur spectral information—color-blind measurements—, recovering standard antibunching from a two-level system. These features are more clearly observed in Fig. 6(a,b).
Similarly to Sec. III [c.f. Fig. 2(c)], the two-sensor CFI, $F_{\Delta}$, inherits a non-trivial spectral structure, shown in the upper-half of the maps shown in the third row in Fig. 5. However, as was the case with the single-sensor CFI and the fluorescence spectrum, it does not directly mirror $g^{(2)}_{\Gamma}$ due to the inherently nonlinear dependence of $F_{\theta}$ on higher-order field moments. The joint probability distribution $p(n_{1},n_{2}|\theta)$ is given by a sum of nonlinear functions of the normally ordered moments of both sensors (see Appendix H),
@@
where we have defined $\hat{D}(\alpha_{1},\alpha_{2})\equiv\hat{D}(\alpha_{1})\hat{D}(\alpha_{2})$, with $\hat{D}(\alpha_{i})=\exp[(\alpha_{i}\hat{\xi}^{\dagger}_{i}-\alpha_{i}^{*}\hat{\xi}_{i})]$, since now each sensor can be independently displaced. While a direct mapping between $g^{(2)}_{\Gamma}$ and $F_{\theta}^{\alpha_{1},\alpha_{2}}$ is not possible, the two-photon correlation function still offers physical insight into the structure of quantum correlations relevant for metrology.
@@
Physically, this situation may occur when, for instance, a path-length mismatch in a the HBT setup [see Fig. 1(a)] introduces time delays large enough to erase any information about photon-photon correlations.
In Fig. 5, the top panels show $F_{\Delta}^{I}$ (blue dot-dashed) as a function of $\Delta_{1}$ for a fixed $\Delta_{2}=2\Omega$, while the lower-half panels in the third row display the ratio $F_{\theta}/F_{\theta}^{I}$, which quantifies the metrological gain from retaining cross-correlations compared to independent measurements. When $F_{\theta}/F_{\theta}^{I}>1$ (red regions), we have a metrological advantage, reaching up to five orders of magnitude near leapfrog resonances, as it is shown in the top panels via the blue dot-dashed curves ($F_{\Delta}^{I}$). For instance, in the aforementioned region, $(\Delta_{1},\Delta_{2})\approx(-2\Omega+\delta\omega,+2\Omega)$, we find $F_{\theta}\approx 10^{5}F_{\theta}^{I}$. Generally, retaining cross-correlations ($F_{\Delta}$) outperforms the independent measurements ($F_{\Delta}^{I}$), as long as spectral resolution is guaranteed. In the case of broadband filtering, the spectral information is washed out, resulting in $F_{\Delta}\approx F_{\Delta}^{I}$ (white regions).
These features are further illustrated in Fig. 6. In panel (a), both $F_{\Delta}$ (blue solid) and $F_{\Delta}^{I}$ (blue dot-dashed) are plotted as a function of $\Gamma$ at a representative leapfrog point $(\Delta_{1},\Delta_{2})=(-2\Omega+\delta\omega,2\Omega)$, while panel (c) displays the ratio $F_{\Delta}/F_{\Delta}^{I}$ across $\Delta_{1}$ and $\Gamma$, clearly highlighting the spectral window where quantum correlations yield maximal metrological benefit.
Figure 6: Impact of sensor linewidth $\Gamma$ on the performance of the estimation protocol for two sensors. (a) $I_{\Delta}$ (grey solid), $F_{\Delta}$ (blue solid), $F_{\Delta}^{I}$ (blue dot-dashed), $F_{\Delta}^{\vec{\alpha}_{\text{opt}}}$ (blue dashed), and $g^{(2)}_{\Gamma}$ (red solid) in terms of $\Gamma$, for a fixed value of the sensors $(\Delta_{1},\Delta_{2})=(-2\Omega+\delta\omega,2\Omega)$. The blue shaded region represents the optimal detection window. (b) $g^{(2)}_{\Gamma}$ and (c) $F_{\Delta}/F_{\Delta}^{I}$ in terms of the frequency of one sensor $\Delta_{1}$ and the sensor linewidth $\Gamma$, while having fixed $\Delta_{2}=2\Omega$. Parameters: $\Delta=0,\ \Omega=10\gamma,\ \Delta_{2}=2\Omega,\ \delta\omega=0.5\Omega,\ \varepsilon=0.5$.
@@
Finally, we note that the mean-field engineering method—introduced earlier in Fig. 3(a) and in Fig. 4—extends naturally to the two-mode scenario. In this case, we must find the optimal displacement for each mode, $F_{\theta}^{\vec{\alpha}_{\text{opt}}}$, where $\vec{\alpha}_{\mathrm{opt}}=(\alpha_{1}^{\mathrm{opt}},\alpha_{2}^{\mathrm{opt}})$ denotes the optimal set of displacements. This definition generalizes straightforwardly to the N-sensor case, where the optimization is carried out over the full set of $N$ coherent displacements. The resulting optimized two-mode CFI is shown as blue dashed curves in the upper panels of Fig. 5 and in Fig. 6(a). In both cases, $F_{\theta}^{\vec{\alpha}_{\mathrm{opt}}}$ exactly saturates the QFI, proving that mean-field engineering provides a general and effective strategy for optimizing the sensitivity of parameter estimation from displaced photon-counting measurements.
V Generality of frequency-resolved quantum metrology
Throughout the manuscript, we focused on the simplest nontrivial quantum-optical system—a coherently driven two-level system—to clearly show the underlying mechanisms and the rich phenomenology enabled by multi-mode frequency filtering and mean-field engineering for quantum metrology. However, the framework we developed is fully general.
To illustrate this broad scope, we now briefly show how the same techniques apply to other relevant physical settings. As an example, we consider the estimation of a nonlinearity $\chi$: first in a single anharmonic oscillator—such as an optical cavity with Kerr nonlinearity or a transmon qubit—and then in a configuration with richer spectral features, where the same system is optomechanically coupled to a resonator yielding Raman sidebands.
V.1 Example 1: Nonlinear harmonic oscillator (transmon qubit)
In this first example, we consider a nonlinear harmonic oscillator (with annihilation operator $\hat{a}$) characterized by a resonant frequency $\omega_{0}$ and a Kerr-type anharmonicity $\chi$, that is coherently driven by a field with amplitude $\Omega$ and frequency $\omega_{L}$. A specific physical setup described by this model is a superconducting transmon qubit coupled to a one-dimensional waveguide [41, 106, 64] [see Fig. 7(a)]. A circulator routes the incoming coherent drive towards the transmon, while directing the outgoing radiation to the detection stage. In the rotating frame of the coherent drive, and under the rotating wave approximation, the Hamiltonian of the driven transmon reads
$\hat{H}_{\mathrm{T}}=\Delta_{a}\hat{a}^{\dagger}\hat{a}-\chi\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+\Omega(\hat{a}+\hat{a}^{\dagger}),$ (24)
where $\Delta_{a}\equiv\omega_{a}-\omega_{L}$ is the transmon-laser detuning. The anharmonicity $\chi$ quantifies the deviation of the transmon from an ideal harmonic oscillator. In the limit $\chi\to 0$, the system behaves as a linear resonator, while for $\chi\to\infty$ the higher excited states become strongly detuned ($\Delta E_{n+1,n}=\omega_{0}-2\chi n$), such that the dynamics effectively reduces to that of a two-level system. Physically, $\chi$ is directly related to the charging energy of the transmon, $E_{C}$, through $\chi\simeq-E_{C}$ [106, 64].
In the Markovian regime, the dissipative dynamics of the system—described by its reduced density matrix $\hat{\rho}_{\mathrm{T}}$—, is governed by the following master equation
$\frac{d\hat{\rho}_{\mathrm{T}}}{dt}=-i[\hat{H}_{\mathrm{T}},\hat{\rho}_{\mathrm{T}}]+\frac{\kappa}{2}\mathcal{D}[\hat{a}]\hat{\rho}_{\mathrm{T}},$ (25)
where $\kappa$ is the relaxation rate of transmon into the waveguide.
Figure 7: Frequency-resolved Fisher information for one-sensor for the estimation of $\theta=\chi$. (a) Sketch of the model: a transmon qubit is coupled to a waveguide in a reflection setup using a circulator. The coherent drive enters through the waveguide and excites the transmon, while the outgoing field is redirected toward the measurement port. The emitted radiation is subsequently frequency-filtered. (b, c) Impact of the sensor linewidth $\Gamma$ on the performance of the estimation protocol for a single sensor. (b) $I_{\chi}$ (grey solid), $F_{\chi}$ (blue solid), $F_{\chi}^{\text{opt}}$ (blue dashed), and $\text{SNR}_{\chi}$ (red solid) as a function of $\Gamma$, shown for a fixed non-linearity $\chi=10\kappa$. (c) $F_{\chi}/\text{SNR}_{\chi}$ as a function of $\chi$ and $\Gamma$. Parameters: $\Delta_{a}=0$, $\Omega=0.1\kappa$, $\varepsilon=1$, $\Delta_{\xi}=2\Omega$.
To analyze the metrological properties of the emitted radiation, we extend Eq. (25) to model the cascaded system of transmon and ancillary sensor. The joint system is described by the density matrix $\hat{\rho}_{\mathrm{T,1}}$, and the corresponding master equation is given by
[…]
Figure 7(b) shows the CFI (blue solid) as a function of the sensor linewidth $\Gamma$ for a particular choice of the nonlinearity, $\chi=10\kappa$. As in the driven-TLS case [see Fig. 3], the CFI exhibits an optimal detection window, here located approximately in the range $\Gamma_{\mathrm{opt}}\approx(10^{-1}\kappa,\,10\,\kappa)$, whereas in both the narrowband and broadband limits the CFI becomes suppressed.
In the narrowband regime $(\Gamma\lesssim 10^{-1}\kappa)$, the CFI collapses onto the SNR (red solid), $F_{\chi}\approx\mathrm{SNR}_{\chi}$, indicating that the estimation is dominated by mean values. Remarkably—and in contrast to the driven-TLS case [see Fig. 3(a)]—both quantities (CFI and SNR) saturate the QFI (grey solid). This proves that, in the narrowband sensor limit, single-detector measurements are already optimal for estimating the transmon nonlinearity $\chi$. In the opposite broadband limit ($\Gamma\gtrsim 10\kappa$), the CFI approaches the SNR as $\Gamma\to\infty$ (not shown), as expected for color-blind detection, where spectral resolution is lost. The QFI remains greater than the CFI for all $\Gamma$, including in the broadband regime, implying the existence of a POVM that outperforms direct photon-counting. By applying mean-field engineering prior to detection [see Fig. 4(a)], we obtain an optimized CFI that saturates the QFI across the full range of $\Gamma$, such that $I_{\chi}\simeq F_{\chi}^{\alpha_{\mathrm{opt}}}$ (blue dashed). This result provides an additional validation of the mean-field engineering approach, showing that even outside the optimal detection region, the estimation sensitivity can still be significantly enhanced.
The above analysis focused on a specific value of the transmon nonlinearity. A similar picture emerges for other values of $\chi$, as shown in Fig. 7(c), which displays the ratio $F_{\chi}/\mathrm{SNR}\chi$ as a function of $\chi$ and $\Gamma$. Regions where this ratio exceeds unity (in red) identify spectral windows in which exploiting the full photon-number distribution provides a metrological advantage over relying only on the mean value, while white regions mark regimes where the field is essentially coherent. Generally, the same behavior reported in panel (b) is present here: a narrowband region where $F\chi\approx\mathrm{SNR}\chi$, an optimal detection window where $F\chi\gg\mathrm{SNR}\chi$, and a broadband limit in which $F\chi\approx\mathrm{SNR}\chi\rightarrow 0$. This shows that the metrological structure of the emitted radiation, when used to estimate the anharmonicity, exhibits a similar dependence on $\chi$ both in the nearly harmonic regime ($\chi\to 0$) and in the strongly anharmonic, qubit-like regime ($\chi\to\infty$). We emphasize, however, that although the ratio $F_{\chi}/\mathrm{SNR}_{\chi}$ varies smoothly with $\chi$, the absolute values of CFI and SNR can differ significantly across regimes. In addition, we observe a peculiar feature, located near $(\chi,\Gamma)\approx(1,10^{-1})\kappa$ and extending over larger $\Gamma$, where $F_{\chi}\approx\text{SNR}_{\chi}$. Further analysis of this system—and of the particular effect just mentioned—is left for a future work.
V.2 Example 2: Optomechanical system
A natural next step beyond the isolated nonlinear oscillator considered in the transmon example above is to explore systems where such nonlinearities interact with additional degrees of freedom. A paradigmatic and technologically relevant example is provided by cavity optomechanics (OM), where the appearance of Raman sidebands—central to many applications in molecular spectroscopy [107]—highlights the broad applicability of our framework in a platform exhibiting an even richer spectral structure. In this context, we note that spectral measurements in optomechanical platforms have also been explored from a complementary perspective, namely to establish metrological bounds for the estimation of time-varying signals and external forces [3, 108] .
In these platforms, the nonlinear coupling between mechanical resonator and optical mode produces a characteristic emission spectrum composed of Stokes and anti-Stokes sidebands associated with phonon creation and annihilation processes [109]. Frequency-resolved correlations have been studied in these platforms in the case of a general OM system [59], or in specific molecular realizations [60, 61], where a rich landscape of frequency-resolved correlations is unraveled due to the intrinsic non-linearities of the system.
We consider a OM system with a single-mode cavity [see Fig. 8(a)], with annihilation operator $\hat{a}$, frequency $\omega_{a}$ and Kerr nonlinearity $\chi$, which is coherently driven with amplitude $\Omega$ at frequency $\omega_{L}$. This optical mode is nonlinearly coupled to a mechanical resonator (with annihilation operator $\hat{b}$) of frequency $\omega_{b}$ via an optomechanical interaction of strength $g$. In the rotating frame of the coherent drive, and under the rotating wave approximation, the Hamiltonian of the OM system takes the form
$\hat{H}_{\mathrm{OM}}=\Delta_{a}\hat{a}^{\dagger}\hat{a}+\chi\hat{a}^{\dagger}\hat{a}^{\dagger}\hat{a}\hat{a}+\Omega(\hat{a}+\hat{a}^{\dagger})\\ +\omega_{b}\hat{b}^{\dagger}\hat{b}+g\hat{a}^{\dagger}\hat{a}(\hat{b}+\hat{b}^{\dagger}),$ (27)
where $\Delta_{a}\equiv\omega_{a}-\omega_{L}$ is the cavity-laser detuning. Dissipation originates from photon loss at rate $\kappa_{a}$, phonon loss at rate $\kappa_{b}$, and thermal phonon excitations characterized by $n_{\mathrm{th}}=(e^{\omega_{b}/k_{B}T}-1)^{-1}$, where $k_{B}$ is the Boltzmann constant and $T$ is the temperature of the phonon bath. In the Markovian regime, the dissipative dynamics of the OM system—described by its reduced density matrix $\hat{\rho}_{\mathrm{OM}}$—, is governed by the following master equation
$\frac{d\hat{\rho}_{\mathrm{OM}}}{dt}=-i[\hat{H}_{\mathrm{OM}},\hat{\rho}_{\mathrm{OM}}]+\frac{\kappa_{a}}{2}\mathcal{D}[\hat{a}]\hat{\rho}_{\mathrm{OM}}\\ +\frac{(n_{\mathrm{th}}+1)\kappa_{b}}{2}\mathcal{D}[\hat{b}]\hat{\rho}_{\mathrm{OM}}+\frac{n_{\mathrm{th}}\kappa_{b}}{2}\mathcal{D}[\hat{b}^{\dagger}]\hat{\rho}_{\mathrm{OM}}.$ (28)
As we did in the previous example, we extend Eq. (28) by including the description of the ancillary sensor, such that the corresponding master equation of the joint system—described by the density matrix $\hat{\rho}_{\mathrm{OM,1}}$—is given by
[…]
Compared to the driven TLS and the transmon qubit, the optomechanical system contains a substantially larger parameter space, $\theta\in\{\Delta_{a},\chi,\Omega,\omega_{b},g,\kappa_{a},\kappa_{b},T\}$, reflecting the richer underlying dynamics. A full metrological characterization of all these parameters lies far beyond the scope of this section. For illustrative purposes—and in continuity with the previous example—we focus on the estimation of the optical nonlinearity, $\theta=\chi$.
Figure 8: Frequency-resolved Fisher information for one-sensor for the estimation of $\theta=\chi$. (a) Sketch of the model: a mechanical resonator ($\hat{b}$) coupled, via a nonlinear interaction of strength $g$, to an optical cavity mode ($\hat{a}$) that exhibits an intrinsic nonlinearity $\chi$. The cavity is also coherently driven (not shown in the sketch). The output field of the cavity is spectrally filtered at frequency $\Delta_{\xi}$ with resolution $\Gamma$. (b, c) Impact of the sensor linewidth $\Gamma$ on the performance of the estimation protocol for a single sensor. (b) $I_{\chi}$ (grey solid), $F_{\chi}$ (blue solid), $F_{\chi}^{\text{opt}}$ (blue dashed), and $\text{SNR}_{\chi}$ (red solid) as a function of $\Gamma$, shown for a fixed sensor frequency $\Delta\xi=-\omega_{b}/2$. (c) $F_{\eta}/\text{SNR}_{\chi}$ as a function of $\Delta\xi$ and $\Gamma$. Parameters: $\Delta_{a}=0$, $\Omega=0.1\kappa_{a}$, $\chi=10^{2}\kappa_{a}$, $\omega_{b}=5\kappa_{a}$, $\kappa_{b}=10^{-2}\kappa_{a}$, $g=\kappa_{a}$, $n_{\text{th}}=10^{-2}$, $\varepsilon=0.1$.
In close analogy with the driven-TLS and transmon examples, the optomechanical system also exhibits a well-defined optimal detection window in which the CFI is maximized, while both the narrowband and broadband limits force the CFI to collapse onto the SNR, as shown in Fig. 8(b). However, the OM nonlinearities generate a richer spectral structure [59]. This is revealed in Fig. 8(c), which displays the ratio $F_{\chi}/\mathrm{SNR}_{\chi}$ as a function of both $\Delta_{\xi}$ and $\Gamma$. As expected, the ratio tends to unity in the narrowband limit (white regions), and would approach unity again in the deep broadband regime (not shown). Importantly, the best metrologically informative frequencies do not coincide with the physical resonances of the source. In the OM system [see Eq. (27)], Stokes and anti-Stokes resonances occur at integer multiples of the mechanical frequency, $\pm m\omega_{b}$, yet optimal regions for estimating $\chi$ appears at other frequencies, e.g., at $\Delta_{\xi}=-\omega_{b}/2$ [as shown in panel (b)]. As it already occurred with the driven-TLS [see Fig. 2(a,c)], this counter-intuitive behavior stems from the fact that the CFI depends on a nontrivial interplay of derivatives of all normally ordered correlators of the filtered field [see Eq. (16)].
Thus, although the scaling of the CFI with the sensor linewidth $\Gamma$ mirrors the behavior observed in the previous cases, the optomechanical setup shows how complex spectral structures—originating from mechanical nonlinearities—can be exploited to enhance parameter estimation.
Our use of a single-sensor configuration and a single parameter in these examples was chosen to provide a simple, self-contained demonstration of how our framework applies to systems that are fundamentally different from the driven TLS on which most of our analysis was focused. A full exploration of the multi-parameter landscape and multi-sensor scenarios in such platforms is beyond the scope of this text. Nevertheless, the results shown here already highlight the potential interest of exploring the metrological potential of frequency-filtered modes in complex quantum-optical systems.
VI Conclusions
We have presented a quantum-metrology framework that harnesses the spectral structure of radiation continuously emitted by open quantum systems. Working with the physical description of the quantum state of the frequency-filtered mode, we used the Fisher information to quantify the precision gained from exploiting the full photon-counting distribution of frequency-filtered modes, as well as the correlations between distinct frequency components revealed through coincidence measurements. Extending this analysis across a range sensor frequencies and linewidths, we identified an optimal filtering regime—where neither spectral nor temporal information is lost—that maximizes sensitivity.